The number of workers n needed to clean a stadium after a game varies inversely as the amount of time t allowed for the cleanup. If it takes 25 workers to clean the stadium when there are 6 hours allowed for the job, how many workers are needed if the stadium must be cleaned in 5 hours?
Answer by stanbon(75887)

he number of workers n needed to clean a stadium after a game varies inversely as the amount of time t allowed for the cleanup.
n = k/t
---------------------
Solve for "k" using "it takes 25 workers to clean the stadium when there are 6 hours allowed for the job"
25 = k/6
k = 150
----
How many workers are needed if the stadium must be cleaned in 5 hours?
n = 150/t
n = 150/5
n = 30 workers
-----------------
Cheers,
Stan H.
-------------